Q. A diffraction grating has 3 x 106 lines per meter. The grating is illuminated by monochromatic plane waves of wavelength 600 nm at normal incidence that forms an interference pattern on a screen located 0.50 m away

a) Find the angle the first bright fringe.

b) Find the angle corresponding to the second bright fringe.

c) Find the angle of the third bright fringe.
Answer #1

Solution: Given:- Distance, d=\frac{1}{3\times 10^{5}}m , wavelength, 600nm .

For first bright fringe the order is given by n=1.

(i) The angle of diffraction normal to grating for first bright fringe is given by:

sin\theta = \frac{n\times \lambda}{d}

sin\theta = \frac{1\times 600\times 10^{-9}}{\frac{1}{3\times 10^{5}}}

sin\theta = 3\times 600\times 10^{-4}

sin\theta = 0.18

\theta = 10.37^{o}

(ii) The angle of diffraction normal to grating for second bright fringe is given by:

Here, n=2.

sin\theta = \frac{n\times \lambda}{d}

sin\theta = \frac{2\times 600\times 10^{-9}}{\frac{1}{3\times 10^{5}}}

sin\theta = 6\times 600\times 10^{-4}

sin\theta = 0.36

\theta = 21.1^{o}

(iii) The angle of diffraction normal to grating for third bright fringe is given by:

Here, n=3.

sin\theta = \frac{n\times \lambda}{d}

sin\theta = \frac{3\times 600\times 10^{-9}}{\frac{1}{3\times 10^{5}}}

sin\theta = 9\times 600\times 10^{-4}

sin\theta = 0.54

\theta = 32.68^{o}